# DriveIndia: An Object Detection Dataset for Diverse Indian Traffic Scenes

### v3

@@ I Introduction
Night-time Data No No No No Yes
@@ IV-B Quantitative Results
Table LABEL:tab:yolo_results summarizes the results for all evaluated models. The YOLO family remains highly competitive, with YOLOv8 achieving the highest $mAP_{50}$ (78.7%), while YOLOv9 offers the best precision/recall balance. Figure  4 visualizes typical DriveIndia scenes by the YOLOv8 Model, reflecting the dataset’s real-world diversity.
@@ IV-C Per-Class Performance Analysis
A detailed breakdown of class-wise metrics is presented in Table II. While common objects like car, truck, and bus achieve high $mAP_{50}$ scores (>0.7>0.7), rare and small-scale classes such as pushcart, pothole, police vehicle, and unmarked speed bump show lower detection accuracy ($mAP_{50{:}95}<0.5$). Figure 5 displays the confusion matrix, highlighting frequent misclassifications, particularly among visually similar or underrepresented classes.
TABLE II: Per-class detection performance on the DriveIndia validation set. Color indicates high, moderate, and low detection scores.
@@ IV-D Discussion: Class-Specific Detection Challenges
Our evaluation [^1] Color categories for Table II For $mAP_{50:95}$, green >>0.50, orange 0.30–0.50, red <<0.30. For $mAP_{50}$, green $\geq$0.75, orange 0.50–0.75, red <<0.50. as seen on Table II shows that common road users—such as car, Auto-rickshaw, truck, and motorcycle—are detected with high accuracy ($mAP_{50}$ above 0.9), reflecting model robustness for prevalent classes in Indian traffic.
@@ V Dataset Availability and Licensing
https://tihan.iith.ac.in/tiand-datasets/

### v4

@@ I Introduction
Night-time Data No No Yes No Yes
@@ IV-B Quantitative Results
Table II summarizes the results for all evaluated models. The YOLO family remains highly competitive, with YOLOv8 achieving the highest $mAP_{50}$ (78.7%), while YOLOv9 offers the best precision/recall balance. Figure  4 visualizes typical DriveIndia scenes by the YOLOv8 Model, reflecting the dataset’s real-world diversity.
TABLE II: Benchmark Results on DriveIndia
Model $\mathbf{mAP_{50}}$ $\mathbf{mAP_{50{:}95}}$ Precision Recall
YOLOv5 0.767 0.545 0.778 0.725
YOLOv8 0.787 0.564 0.775 0.754
YOLOv9 0.779 0.560 0.791 0.734
YOLOv11 0.717 0.506 0.714 0.660
RT-DETR 0.765 0.553 0.764 0.728
EfficientDet 0.726 0.512 0.743 0.678
@@ IV-C Per-Class Performance Analysis
A detailed breakdown of class-wise metrics is presented in Table III. While common objects like car, truck, and bus achieve high $mAP_{50}$ scores (>0.7>0.7), rare and small-scale classes such as pushcart, pothole, police vehicle, and unmarked speed bump show lower detection accuracy ($mAP_{50{:}95}<0.5$). Figure 5 displays the confusion matrix, highlighting frequent misclassifications, particularly among visually similar or underrepresented classes.
TABLE III: Per-class detection performance on the DriveIndia ‘Validation set’. Color indicates high, moderate, and low detection scores.
@@ IV-D Discussion: Class-Specific Detection Challenges
Our evaluation [^1] Color categories for Table III For $mAP_{50:95}$, green >>0.50, orange 0.30–0.50, red <<0.30. For $mAP_{50}$, green $\geq$0.75, orange 0.50–0.75, red <<0.50. as seen on Table III shows that common road users—such as car, Auto-rickshaw, truck, and motorcycle—are detected with high accuracy ($mAP_{50}$ above 0.9), reflecting model robustness for prevalent classes in Indian traffic.
@@ V Dataset Availability and Licensing
https://tihan.iith.ac.in/TiAND.html